The magnetic intensity H at the centre of a long solenoid carrying a current of 2.0 A, is found to be 1500 A/m. Find the number of turns per centimeter of the solenoid.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. 1. B = μ0ni, H = B/μ base 0 ⇒ H = ni ⇒ 1500 A/m = n× 2 ⇒ n = 750 turns/meter ⇒ n = 7.5 turns/cm